OpenStudy (anonymous):

how would you prepare 5000gramsof 0.5 molal sodium hydroxide solution

OpenStudy (cwrw238):

1 liter 0.5 molar sodium hydroxide contains 0.5 * 1 mole of the substance

OpenStudy (anonymous):

10ml 5.000 M NaOH solution + 90ml H2O. Explanation: 5.000 divided by 0.5000 = 10. So you want it 1/10 of the strength it is now. Thus, stick 1/10 of the final volume with 9/10 of the final volume. You now have gone from a 5 M solution to a 0.5 M solution.

OpenStudy (anonymous):

MV = MV 5M * V = 100ml * 0.5M V = 100ml * 0.5M/5M = 10ml

OpenStudy (zpupster):

1)Determine grams of water in 1000 grams solution In 0.5 molal there is .5 mole NaOH in 1000 grams solvent 0.5 mole NaOH X (40 grams / 1 mole) = 20 grams NaOH 2)Total grams solution = grams of NaOH + grams of solvent 1000 grams solution = 20 grams + grams solvent grams of solvent = 1000 - 20 = 980 grams solvent per 1000 grams solution 5000 grams solution X 980 grams water / 1000 grams solution = 4900 grams water grams NaOH = Total grams of solution - grams of solvent grams NaOH = 5000 - 4900 = 100 grams NaOH